3 2 1

Polarizers 1 and 3 have their axes of polarization indicated by the black solid lines, perpendicular to each other. Polarizer 2 has an axis of polarization that has an angle of ? = 54.1° with respect to the polarization axis of polarizer 1. What percentage of the initial (unpolarized) light intensity passes through?

A: 11.28 B: 15.00 C: 19.95 D: 26.54 E: 35.30 F: 46.95 G: 62.44 H: 83.04
Answer #1

When light passes through first polarizer, it's intensity will be reduced to half

I1 = I0/2

After that when light passes thorugh 2nd polarizer, Intensity will be

I2 = I1*(cos A1)^2

After that when light passes thorugh 3rd polarizer, Intensity will be

I3 = I2*(cos A2)^2

I3 = I1*(cos A1)^2*(cos A2)^2

I3 = (I0/2)*(cos A1)^2*(cos A2)^2

I3/I0 = (1/2)*(cos A1)^2*(cos A2)^2

Now using given values:

A1 = 54.1 deg

A2 = 90 - 54.1 = 35.9 deg

I3/I0 = (1/2)*(cos 54.1 deg)^2*(cos 35.9 deg)^2

I3/I0 = 0.1128

In percentage = 11.28%

Correct option is A

Please Upvote.